OpenStudy (anonymous):

how to find the simplified form of an expression. Give your answer in scientific notation (8* 10^7)(7*10^4)

OpenStudy (jdoe0001):

\(\large \bf (8\times 10^{\color{blue}{ 7}})(7\times 10^{\color{blue}{ 4}})\implies (8\times 7)\times 10^{{\color{blue}{ 7+4}}}\)

OpenStudy (anonymous):

Thanks c:

OpenStudy (jdoe0001):

yw